# Reducing Hallucinations in Language Model-based SPARQL Query Generation Using Post-Generation Memory Retrieval

### v2

@@ 7 Conclusion
In this work, we propose PGMR, a novel SPARQL query generation framework, which incorporates a non-parametric memory module to address the issue of hallucinations in LLM-generated queries. In this framework, LLMs are responsible for generating intermediate SPARQL queries where KG elements like URIs are represented in natural language and enclosed in special tokens, instead of directly producing full SPARQL queries. A retriever with a non-parametric memory module is utilized to seamlessly integrate the retrieved URIs into the intermediate query, resulting in the final SPARQL query. Despite its simplicity, PGMR consistently demonstrates strong performance in SPARQL query generation across a range of datasets, data distribution shifts, and LLMs in both finetuned and few-shot settings. Remarkably, PGMR leads to a considerable decrease in URI hallucinations, achieving a near-complete elimination of this problem across various instances.
@@
Even though PGMR exhibits markedly improved performance in generating correct URIs and reducing hallucinations, its effectiveness is tempered by the LLM’s ability to construct SPARQL query structures accurately. This is especially evident when looking at the results on the QALD-10 dataset (see Table 3), which features notably complex queries involving diverse clauses and literal expressions compared to LCQUAD 2.0. This limitation may be addressed by using more powerful large LLMs pre-trained specifically over SPARQL query data, combined with the PGMR generation paradigm, but we leave that for future work.
@@ Acknowledgements
We thank NSERC and Samsung for supporting this work and CIFAR for their support under the Canada CIFAR AI Chair program.
Appendix A Appendix
A.1 Hyperparameters
Our T5-small model has 60,506,624 trainable parameters and was finetuned on a single 80GB Nvidia A100 GPU within 14 hours with a batch size of 128. We used a learning rate of 0.0015 for finetuning T5 and train for 150 epochs on LCQUAD 2.0 and 30 epochs on QALD-10. Owing to the limited training data for QALD-10, we finetuned our versions of T5 pre-trained on the much larger LCQUAD 2.0 dataset.
For our experiments with Llama 3.1 8B, we fine-tuned the Llama-3.1-8B-Instruct model, training it for 10 epochs using the AdamW optimizer with a learning rate 0.0001, no weight decay, and a fixed seed 42 to ensure reproducibility. Mixed precision training and quantization were enabled to improve efficiency. We employed a parameter-efficient finetuning strategy using LoRA, with a low-rank dimension of 8, a scaling factor (alpha) of 32, and a dropout rate of 0.05. The complete training process took 2 hours, 20 minutes, and 49 seconds, and notably, the model achieved convergence at epoch 5, underscoring its rapid convergence and effectiveness in capturing key aspects of the target data.

### v3

@@ 8 Conclusion
We presented PGMR (Post-Generation Memory Retrieval), a modular framework that addresses URI hallucination in LLM-generated SPARQL queries by separating query logic generation from retrieval. LLMs generate intermediate queries with URI placeholders and associated URI natural language mappings, which a non-parametric retriever then resolves into accurate URIs. Despite its simplicity, PGMR consistently achieved superior semantic query match (SQM) and answer-level F1 scores across diverse LLMs and datasets, yielding a near-complete elimination of URI hallucinations. Crucially, the system demonstrates exceptional robustness, both in its native ability to refuse to answer queries based on confidence thresholds and its resilience to memory scaling, maintaining high SQM with only minimal performance degradation even when memory was expanded nine-fold with distracting entities.
@@
Even though PGMR exhibits markedly improved performance in generating correct URIs and reducing hallucinations, its effectiveness is tempered by the LLM’s ability to construct SPARQL query structures accurately. This is especially evident when looking at the results on the QALD-10 dataset (see Table 3), which features notably complex queries involving diverse clauses and literal expressions compared to LCQUAD 2.0. This limitation may be addressed by using more powerful, large LLMs pre-trained specifically on SPARQL query data, combined with the PGMR generation paradigm, or by combining the advantages of RAG and PGMR, adding a retrieval step both before and after generation. However, we leave that for future work.
@@ Acknowledgements
We thank NSERC and Samsung for supporting this work and CIFAR for their support under the Canada CIFAR AI Chair program. Additionally, we thank IVADO, as this project was undertaken thanks to funding from IVADO and the Canada First Research Excellence Fund. We also thank Luis Lara for his help with running some of the experiments in the early part of this work.
Appendix A Appendix A
Figure 4: The retriever finds each natural language URI label in the generated intermediate query using the special tokens and retrieves the corresponding URI from memory using similarity search, as discussed in Section 4.2
Figure 5: RAG-based SPARQL generation workflow. A retriever first fetches $k$ URIs based on latent similarity to the input question, which are then utilized by the LLM to synthesize the final query, as discussed in Section A.1
Figure 6: The prompt for RAG includes k retrieved URIs along with the question as input to the LLM. In our study, we set $k$ to 10 in our RAG baseline.
A.1 Retrieval Augmented Generation (RAG)
Retrieval Augmented Generation (RAG) (Khandelwal et al., 2019; Lewis et al., 2020b; Guu et al., 2020; Lewis et al., 2020a) has become a popular technique for injecting external information into LLMs, helping them generate content grounded in an external knowledge base like a KG. As seen in Figure 5, we propose RAG for SPARQL query generation, where the question first goes to a retriever (discussed in further detail in Section 4.2) which returns $k$ URIs from the memory whose labels are the closest to the question text in latent space, where $k$ is a positive integer. The URIs and their labels are appended to the question to create the prompt for the LLM (see Figure 6). The LLM then uses this information to generate the SPARQL query corresponding to the given question.
Appendix B Appendix B
Unlike the evaluation approach utilized by Banerjee et al. (2022) and Qi et al. (2024), which involves generating multiple queries for each question and selecting the first query that successfully executes and returns an answer as the predicted query, our evaluation generates only a single query from our models.
B.1 Hyperparameters
Our T5-small model has 60,506,624 trainable parameters and was finetuned on a single 80GB Nvidia A100 GPU within 14 hours with a batch size of 128. We used a learning rate of 0.0015 for finetuning T5 and trained for 150 epochs on LCQUAD 2.0 and 30 epochs on QALD-10. Due to the limited training data for QALD-10, we fine-tuned our versions of T5, which were pre-trained on the much larger LCQUAD 2.0 dataset.
For our experiments with Llama 3.1 8B, we fine-tuned the Llama-3.1-8B-Instruct model, training it for 10 epochs using the AdamW optimizer with a learning rate of 0.0001, no weight decay, and a fixed seed of 42 to ensure reproducibility. Mixed precision training and quantization were enabled to improve efficiency. We employed a parameter-efficient finetuning strategy using LoRA, with a low-rank dimension of 64, a scaling factor (alpha) of 128, and a dropout rate of 0.05. The complete training process took around 3 hours and 30 minutes. Notably, the model achieved convergence at epoch 5, underscoring its rapid convergence and effectiveness in capturing key aspects of the target data.